# Parametric MMD Estimation with Missing Values: Robustness to Missingness and Data Model Misspecification

### v1

@@ 1 Introduction
A promising alternative to MLE is Maximum Mean Discrepancy (MMD) estimation, which selects parameters by minimizing the MMD between the postulated and observed distributions, rather than relying on the Kullback-Leibler divergence as in MLE. This approach is known to exhibit desirable robustness properties (Briol et al., (2019); Chérief-Abdellatif and Alquier, (2022)). Building on the idea of safeguarding against deviations in both the assumed data distribution and the missingness mechanism, we extend MMD estimation to the case of missing values. Our first step is to reformulate the MMD estimator as an M-estimator, which allows us to adapt it to the more challenging setting of missing data. We then establish that the resulting estimator is consistent and asymptotically normal under MCAR, while remaining robust to certain deviations from both the assumed data model and the missingness mechanism. Specifically, we show that under regularity conditions, the MMD estimator converges to the parameter associated with the closest distribution (in MMD) to the true conditional within the model class. We derive explicit bounds on this MMD distance, which split into two components: one corresponding to errors from model misspecification and the other accounting for deviations from the MCAR assumption in the missingness mechanism. To facilitate practical implementation, we extend an existing stochastic gradient descent (SGD) algorithm for MMD estimation to handle missing values. This allows for a straightforward adaptation of the algorithm to a wide range of parametric models, maintaining its generality, flexibility, and computational efficiency while providing rigorous robustness guarantees.
The paper is organized as follows. After discussing related work and notation, we first give some background on MMD and maximum likelihood estimation and motivate our approach in Section 2. Section 3 then presents our estimator and its asymptotic properties. Section 4 discusses the general robustness guarantees of the new estimator. Finally, Section 5 provides several examples.
@@ 1.1 Related Work
The central statistical tool in our framework is the Maximum Mean Discrepancy (MMD) Gretton et al., (2007, 2012), a distance measure that has gained increasing attention in parametric and nonparametric estimation. Initially introduced for hypothesis testing, MMD has since been established as a general estimation principle by Briol et al., (2019) and Chérief-Abdellatif and Alquier, (2022), who showed that MMD-based estimators are consistent, robust to model misspecification, and computationally tractable. These properties make MMD particularly well-suited for scenarios involving complex data distributions and deviations from assumed models. The application of MMD estimation has been extended to various settings, including generalized Bayesian inference Chérief-Abdellatif and Alquier, (2020); Pacchiardi et al., (2024), Approximate Bayesian Computation (ABC) Park et al., (2016); Mitrovic et al., (2016); Kajihara et al., (2018); Bharti et al., (2021); Legramanti et al., (2025), Bayesian nonparametric learning Dellaporta et al., (2022), generative adversarial networks Dziugaite et al., (2015); Li et al., (2015); Sutherland et al., (2016); Li et al., (2017); Bińkowski et al., (2018), quantization Teymur et al., (2021), copula estimation Alquier et al., (2023), label shift quantification Zhang et al., (2013); Iyer et al., (2014); Dussap et al., (2023), and regression Alquier and Gerber, 2024b , among others.
Our work extends this line of research by applying MMD-based estimation to the context of missing data, demonstrating that it provides a natural and robust alternative to traditional methods while avoiding explicit specification of the missingness mechanism.
@@ 2.2 MMD estimation with no missing data
and the KL divergence is not bounded and does not induce the desirable robustness properties inherent to the MMD. Consider for example the case where the data are realizations of $n$ i.i.d. random variables $X_{1},\dots,X_{n}$ actually drawn from the mixture:
${\mathbb{P}_{X}}=\left(1-\epsilon\right)\cdot\mathcal{N}(\theta^{*},1)+\epsilon\cdot\mathcal{N}(\xi,1),$
where the distribution of interest $\mathcal{N}(\theta^{*},1)$ is contaminated by another univariate normal $\mathcal{N}(\xi,1)$. Then $\theta_{\infty}^{\textnormal{ML}}=(1-\epsilon)\theta^{*}+\epsilon\xi$, which can be quite far from $\theta^{*}$ when $|\epsilon(\xi-\theta^{*})|$ is large when compared to $\theta^{*}$. This illustrates the non-robustness of the MLE to small deviations from the parametric assumption.
@@ 2.3 Maximum Likelihood Estimation under missing data
$\theta_{\infty}^{\textnormal{ML}}={\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\textnormal{KL}\left({\mathbb{P}_{X\mid M}^{(M)}}\|{P^{(M)}_{\theta}}\right)\right],$
@@ 3.2 Computation using SGD
[…]
@@ 3.3 Asymptotic analysis
In this section, we analyze the properties of $\theta_{n}^{\textnormal{MMD}}$, without making any assumption on the nature of the missingness mechanism. The conditions we formulate here are broad and may differ slightly from those typically used in the MMD literature. Existing approaches generally fall into two categories: assumptions on the density Key et al., (2021); Alquier and Gerber, 2024b or on the generator Briol et al., (2019). Some works Oates, (2022); Alquier et al., (2023) have proposed more general conditions that encompass both perspectives, though these can be challenging to establish due to the diverse nature of the underlying random objects.
@@
  • The limit is no longer the minimizer of the MMD distance to the true whole distribution ${\mathbb{P}_{X}}$ as in the complete data setting. It now corresponds to the best average (over $M\sim{\mathbb{P}_{M}}$) approximation of the observed marginal of the conditional ${\mathbb{P}_{X\mid M}}$ in terms of MMD distance, as for the ignoring MLE, and is equal to $\theta^{*}$ as soon as the model is well-specified and the missingness mechanism is MCAR. This setting of the mechanism is crucial, as under M(N)AR, distribution shifts are possible when moving from one missingness pattern to the next. In fact, as discussed in Näf et al., (2024) and others, MAR can be defined as a restriction on the arbitrary distribution shifts that are allowed under MNAR. Intuitively, if these distribution shifts from one pattern to the next are not too extreme $\theta_{\infty}^{\textnormal{MMD}}$ will “reasonably close” to the truth.
@@ 5 Examples
We investigate in this subsection explicit applications of our theory for specific examples of MNAR missingness mechanisms. We provide three of them: the first one involves a truncation mechanism; the second one is a Huber contamination setting for the missingness mechanism as explored in Ma et al., (2024); while the third one allows for adversarial contamination of the missingness mechanism.
@@ 5.1 MNAR Truncation Mechanism
A one-dimensional dataset $\{X_{i}\}_{i}$ composed of i.i.d. copies of $X$ is collected, but we only observe a subsample of it $\{X_{i}:M_{i}=0\}_{i}$, where $M_{i}$ is an observed missingness random variable equal to $1$ if $X_{i}$ is observed, and 00 otherwise. The true missingness mechanism is actually MNAR, with $X$ being observed only when it belongs to some large subset $\mathcal{I}$ of $\mathbb{R}$:
@@ 5.2 Huber’s Contamination Mechanism
An application of this result in the univariate Gaussian model would lead to a rate of order $\max(\varepsilon/\alpha(1-\varepsilon),\{n\alpha(1-\varepsilon)\}^{-1/2})$ (for $|\theta_{n}^{\textnormal{MMD}}-\theta^{*}|$) when $\varepsilon=\epsilon$, which aligns with the minimax optimal (high-probability) rate established in Table 1 of Ma et al., (2024). The two contamination settings are however different, since we are interested here in the separate contamination of the complete data marginal distribution and of the missingness mechanism, while Ma et al., (2024) rather focus on the contamination of the joint distribution. Nevertheless, while our contamination model for $\epsilon=\varepsilon$ does not directly correspond to the so-called arbitrary contamination setting of Ma et al., (2024), our framework coincides with their realizable contamination scenario when $\epsilon=0$, meaning that ${\mathbb{P}_{X}}={P_{\theta^{*}}}$. Interestingly, the authors highlighted a surprising result in this case - and which seems to be unique to the Gaussian model; they demonstrated that consistency towards $\theta^{*}$ can still be achieved as $n\to\infty$, even if $\varepsilon>0$. They also provided a simple estimator - the average of the extreme values - that satisfies such consistency. While this phenomenon is indeed surprising, it comes at a cost: the convergence is logarithmic in the sample size, specifically ${\log(n\alpha(1-\varepsilon))}^{-1/2}$. While our MMD estimator does not achieve consistency towards $\theta^{*}$, its convergence rate towards $\theta_{\infty}^{\textnormal{MMD}}$ is in $\{(n\alpha(1-\varepsilon))\}^{-1/2}$, thus outperforming the average of extremes estimator in this regard for finite samples. Figure 1 illustrates this behavior for the two estimators in a Gaussian mean example: while the MMD quickly suffers an error of $\varepsilon=0.1$ as $n$ increases, the average of extremes moves very slowly to the true value $\theta^{*}=0$.
@@ 5.3 Adversarial Contamination Mechanism
Let us now introduce a more sophisticated adversarial contamination setting for the missingness mechanism in the one-dimensional case: from the dataset $\{X_{i}\}_{i}$ composed of i.i.d. copies of $X\sim{\mathbb{P}_{X}}$, only a subsample of it $\{X_{i}:M_{i}=0\}_{i}$ is observed according to a MCAR missingness mechanism $\mathbb{P}[M=0|X=x]=\alpha$ for any $x$. However, the selected observed examples $M_{i}$’s have been modified by an adversary, and we finally observe $\{X_{i}:\widetilde{M}_{i}=0\}_{i}$ where at most |$|\{i:M_{i}=0\}|\cdot\varepsilon/\alpha$ (with $0<\varepsilon<\alpha$) missing variables $\widetilde{M}_{i}$ are different from the original $M_{i}$.
The choice |$|\{i:M_{i}=0\}|\cdot\varepsilon/\alpha$ with $\varepsilon<\alpha$ is such that the adversary cannot remove all the examples in the finally observed dataset by switching all $M_{i}$’s equal to 00 to the value $\widetilde{M}_{i}=1$, and such that there are on average (over the sample $\{M_{i}\}_{i}$) at most $n\varepsilon$ contaminated values (recall that contamination is in the missing variable $M$). The following theorem provides a bound that is valid for any finite value of the sample size $n$:
@@
Consequences of this subtle difference can be important in the study of minimax theory for population mean estimation problems, with completely different minimax rates when compared to Huber’s setting. It is very unlikely that consistency towards $\theta^{*}$ is possible any longer in the realizable setting where ${\mathbb{P}_{X}}={P_{\theta^{*}}}$, and very likely that optimal estimators for Huber’s model behave extremely bad in this adversarial model. For instance, the average of extreme procedure becomes a terrible estimator under adversarial contamination of the mechanism: it is sufficient for the adversary to remove the $\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|-1$ smallest points and observe the largest one for a symmetric distribution to mislead the average of extremes estimator, no matter the value of the contamination ratio. At the opposite, the MMD estimator still behaves correctly in this setting, with an error which is (almost) linear in $\varepsilon$. Figure 1 illustrates this behavior: while the MMD is stable around $\varepsilon=0.1$, the average of extremes moves farther away from the true value of the mean as $n$ increases. The minimax optimality in the adversarial contamination setting is an open question.
@@
Figure 1: MLE, MMD (with Gaussian kernel) and Average of Extremes Estimator for a Gaussian mean estimation problem following Example 5.2 (Huber contamination) in light blue and Example 5.3 (Adversarial Contamination) in red. The shaded regions correspond to the empirical 50% quantile range obtained by rerunning the experiment $1000$ times for each $n$. In both settings, we simulate $X\sim N(0,1)$ without contamination. For the case of Huber contamination, the MCAR mechanism with $\alpha=0.5$ probability of missingness is contaminated by an MNAR mechanism for which $M=0$ iff $X>0$. For the adversarial mechanism, each $M_{i}$ is first generated following an MCAR mechanism with $\alpha=0.5$ probability of missingness, and in a second step, the $M_{i}$’s corresponding to the $12\varepsilon\cdot|\{i:M_{i}=0\}|-$ 1 smallest values are set to 1, while the $M_{i}$ corresponding to the largest value is set to 00. In both cases, $\epsilon=0.1$. Implementation was performed in R with the help of the R package regMMD (Alquier and Gerber, 2024a, ).

### v2

@@ 1 Introduction
A promising alternative to MLE is Maximum Mean Discrepancy (MMD) estimation, which selects parameters by minimizing the MMD between the postulated and observed distributions, rather than relying on the Kullback-Leibler divergence as in MLE. This approach is known to exhibit desirable robustness properties (Briol et al., (2019); Chérief-Abdellatif and Alquier, (2022)). Building on the idea of safeguarding against deviations in both the assumed data distribution and the missingness mechanism, we extend MMD estimation to the case of missing values. Our first step is to reformulate the MMD estimator as an M-estimator, which allows us to adapt it to the more challenging setting of missing data. We then establish that the resulting estimator is consistent and asymptotically normal under MCAR, while remaining robust to certain deviations from both the assumed data model and the missingness mechanism. Specifically, we show that under regularity conditions, the MMD estimator converges to the parameter associated with the closest distribution (in MMD) to the true conditional within the model class. We derive explicit bounds on this MMD distance, which split into two components: one corresponding to errors from model misspecification and the other accounting for deviations from the MCAR assumption in the missingness mechanism. To facilitate practical implementation, we extend an existing stochastic gradient descent (SGD) algorithm for MMD estimation to handle missing values. This allows for a straightforward adaptation of the algorithm to a wide range of parametric models, maintaining its generality, flexibility, and computational efficiency while providing rigorous robustness guarantees. In the special case of the Gaussian distribution with fixed covariance matrix, this leads to a robust estimator of the mean under missing values. While estimation of a Gaussian mean under contamination has been the subject of active research in the last decade with complete data (see e.g., Diakonikolas and Kane, (2019); Loh, (2024); Anderson and Phillips, (2025) and the references therein for an overview), much less attention has been paid to the problem with missing data. Two recent exceptions are Hu and Reingold, (2021) who study the problem theoretically in great detail and Ma et al., (2024) who improve on their results.
The paper is organized as follows. After discussing related work and notation, we first give some background on MMD and maximum likelihood estimation and motivate our approach in Section 2. Section 3 then presents our estimator and its asymptotic properties. Section 4 discusses the general robustness guarantees of the new estimator. Finally, Sections 5 and 6 provide several examples. Code to replicate the examples can be found on https://github.com/JeffNaef/MissingnessMMD.
@@ 1.1 Related Work
The central statistical tool in our framework is the Maximum Mean Discrepancy (MMD) (Gretton et al., , 2007; Gretton et al., 2012a, ), a distance measure that has gained increasing attention in parametric and nonparametric estimation. Initially introduced for hypothesis testing, MMD has since been established as a general estimation principle by Briol et al., (2019) and Chérief-Abdellatif and Alquier, (2022), who showed that MMD-based estimators are consistent, robust to model misspecification, and computationally tractable. These properties make MMD particularly well-suited for scenarios involving complex data distributions and deviations from assumed models. The application of MMD estimation has been extended to various settings, including generalized Bayesian inference (Chérief-Abdellatif and Alquier, , 2020; Pacchiardi et al., , 2024; Frazier et al., , 2024; Shen et al., , 2024), Approximate Bayesian Computation (ABC) (Park et al., , 2016; Mitrovic et al., , 2016; Kajihara et al., , 2018; Bharti et al., , 2021; Legramanti et al., , 2025), Bayesian nonparametric learning (Dellaporta et al., , 2022), generative adversarial networks (Dziugaite et al., , 2015; Li et al., , 2015; Sutherland et al., , 2016; Li et al., , 2017; Bińkowski et al., , 2018), quantization (Teymur et al., , 2021), copula estimation (Alquier et al., , 2023), label shift quantification (Zhang et al., , 2013; Iyer et al., , 2014; Dussap et al., , 2023), and regression (Alquier and Gerber, 2024b, ), among others.
Our work extends this line of research by applying MMD-based estimation to the context of missing data, demonstrating that it provides a natural and robust alternative to traditional methods while avoiding explicit specification of the missingness mechanism. Although MMD-based hypothesis testing has recently been extended to missing data scenarios (Matabuena et al., , 2023; Zeng et al., , 2024), to the best of our knowledge, this is the first work to address this problem within the framework of robust estimation.
@@ 1.2 Notation
Note that for all the expectations in $M$, the ”empty pattern” is (implicitly) excluded, i.e.
$\mathbb{E}_{M}\left[f(M)\right]=\mathbb{E}_{M}\left[\mathbbm{1}(M\neq\{1\}^{d})f(M)\right]=\sum_{m\neq\{1\}^{d}}p_{m}f(m).$
The indicator will be omitted for ease of presentation.
@@ 2.2 MMD estimation with no missing data
and the KL divergence is not bounded and does not induce the desirable robustness properties inherent to the MMD. Consider for example the case where the data are realizations of $n$ i.i.d. random variables $X_{1},\dots,X_{n}$ actually drawn from the mixture: ${\mathbb{P}_{X}}=\left(1-\epsilon\right)\cdot\mathcal{N}(\theta^{*},1)+\epsilon\cdot\mathcal{N}(\xi,1),$ where the distribution of interest $\mathcal{N}(\theta^{*},1)$ is contaminated by another univariate normal $\mathcal{N}(\xi,1)$. Then $\theta_{\infty}^{\textnormal{ML}}=(1-\epsilon)\theta^{*}+\epsilon\xi$, which can be quite far from $\theta^{*}$ when $|\epsilon(\xi-\theta^{*})|$ is large when compared to $\theta^{*}$. This illustrates the non-robustness of the MLE to small deviations from the parametric assumption. An empirical example illustrating this is given in Section 6.
@@ 2.3 Maximum Likelihood Estimation under missing data
$\theta_{\infty}^{\textnormal{ML}}=\arg\min_{\theta\in\Theta}{\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\textnormal{KL}\left({\mathbb{P}_{X\mid M}^{(M)}}\|{P^{(M)}_{\theta}}\right)\right$],
@@ 3.2 Computation using SGD
Assume that each $P_{\theta}$ has a density $p_{\theta}$ w.r.t. Lebesgue’s measure. Assume that for any $x$, the map $\theta\mapsto p_{\theta}(x)$ is differentiable w.r.t. $\theta$, and there exists a nonnegative function $g(x,x^{\prime})$ such that for any $\theta\in\Theta$, $\left|k(x,x^{\prime})\nabla_{\theta}[p_{\theta}(x)p_{\theta}(x^{\prime})]\right|\leq g(x,x^{\prime})$ and $\int\int g(x,x^{\prime})dxdx^{\prime}<\infty$. Then the gradient of Objective (6) is given by:
@@ 3.3 Asymptotic analysis
In this section, we analyze the properties of $\theta_{n}^{\textnormal{MMD}}$, without making any assumption on the nature of the missingness mechanism. The conditions we formulate here are broad and may differ slightly from those typically used in the MMD literature. Existing approaches generally fall into two categories: assumptions on the density (Key et al., , 2021; Alquier and Gerber, 2024b, ) or on the generator (Briol et al., , 2019). Some works Oates, (2022); Alquier et al., (2023) have proposed more general conditions that encompass both perspectives, though these can be challenging to establish due to the diverse nature of the underlying random objects.
@@
  • The limit is no longer the minimizer of the MMD distance to the true whole distribution $\mathbb{P}_{X}$} as in the complete data setting. It now corresponds to the best average (over $M\sim{\mathbb{P}_{M}}$) approximation of the observed marginal of the conditional $\mathbb{P}_{X\mid M}$} in terms of MMD distance, as for the ignoring MLE, and is equal to $\theta^{*}$ as soon as the model is well-specified and the missingness mechanism is MCAR. This setting of the mechanism is crucial, as under M(N)AR, distribution shifts are possible when moving from one missingness pattern to the next. In fact, as discussed in Molenberghs et al., (2008); Näf et al., (2024) and others, MAR can be defined as a restriction on the arbitrary distribution shifts that are allowed under MNAR. Intuitively, if these distribution shifts from one pattern to the next are not too extreme $\theta_{\infty}^{\textnormal{MMD}}$ will “reasonably close” to the truth.
@@ 5 Examples
We investigate in this subsection explicit applications of our theory for specific examples of MNAR missingness mechanisms. We provide three of them: the first one involves a truncation mechanism; the second one is a Huber contamination setting for the missingness mechanism; while the third one allows for adversarial contamination of the missingness mechanism.
@@ 5.1 MNAR Truncation Mechanism
A one-dimensional dataset $\{X_{i}\}_{i}$ composed of i.i.d. copies of $X$ is collected, but we only observe a subsample of it $\{X_{i}:M_{i}=0\}_{i}$, where $M_{i}$ is an observed missingness random variable equal to 0 if $X_{i}$ is observed, and $1$ otherwise. The true missingness mechanism is actually MNAR, with $X$ being observed only when it belongs to some large subset $\mathcal{I}$ of $\mathbb{R}$:
@@ 5.2 Huber’s Contamination Mechanism
An application of this result in the univariate Gaussian model would lead to a rate of order $\max(\varepsilon/\alpha(1-\varepsilon),\{n\alpha(1-\varepsilon)\}^{-1/2})$ (for $|\theta_{n}^{\textnormal{MMD}}-\theta^{*}|$) when $\varepsilon=\epsilon$, which aligns with the minimax optimal (high-probability) rate established in Table 1 of Ma et al., (2024). The two contamination settings are however different, since we are interested here in the separate contamination of the complete data marginal distribution and of the missingness mechanism, while Ma et al., (2024) rather focus on the contamination of the joint distribution. Nevertheless, while our contamination model for $\epsilon=\varepsilon$ does not directly correspond to the so-called arbitrary contamination setting of Ma et al., (2024), our framework coincides with their realizable contamination scenario when $\epsilon=0$, meaning that ${\mathbb{P}_{X}}={P_{\theta^{*}}}$. Interestingly, the authors highlighted a surprising result in this case - and which seems to be unique to the Gaussian model; they demonstrated that consistency towards $\theta^{*}$ can still be achieved as $n\to\infty$, even if $\varepsilon>0$. They also provided a simple estimator - the average of the extreme values - that satisfies such consistency. This phenomenon is indeed surprising, and it comes at a cost: the convergence is logarithmic in the sample size, specifically ${\log(n\alpha(1-\varepsilon))}^{-1/2}$. While our MMD estimator does not achieve consistency towards $\theta^{*}$, its convergence rate towards $\theta_{\infty}^{\textnormal{MMD}}$ is in $\{(n\alpha(1-\varepsilon))\}^{-1/2}$, thus outperforming the average of extremes estimator in this regard for finite samples. The blue curves in Figure 1 illustrate this behavior for the two estimators in a Gaussian mean example: while the MMD quickly suffers an error of $\varepsilon=0.1$ as $n$ increases, the average of extremes moves very slowly to the true value $\theta^{*}=0$.
@@ 5.3 Adversarial Contamination Mechanism
Let us now introduce a more sophisticated adversarial contamination setting for the missingness mechanism in the one-dimensional case: from the dataset $\{X_{i}\}_{i}$ composed of i.i.d. copies of $X\sim{\mathbb{P}_{X}}$, only a subsample of it $\{X_{i}:M_{i}=0\}_{i}$ is observed according to a MCAR missingness mechanism $\mathbb{P}[M=0|X=x]=\alpha$ for any $x$. However, the selected observed examples $M_{i}$’s have been modified by an adversary, and we finally observe $\{X_{i}:\widetilde{M}_{i}=0\}_{i}$ where at most $|\{i:M_{i}=0\}|\cdot\varepsilon/\alpha$ (with $0<\varepsilon<\alpha$) missing variables $\widetilde{M}_{i}$ are different from the original $M_{i}$.
The choice $|\{i:M_{i}=0\}|\cdot\varepsilon/\alpha$ with $\varepsilon<\alpha$ is such that the adversary cannot remove all the examples in the finally observed dataset by switching all $M_{i}$’s equal to 0 to the value $\widetilde{M}_{i}=1$, and such that there are on average (over the sample $\{M_{i}\}_{i}$) at most $n\varepsilon$ contaminated values (recall that contamination is in the missing variable $M$). The following theorem provides a bound that is valid for any finite value of the sample size $n$:
@@
Consequences of this subtle difference can be important in the study of minimax theory for population mean estimation problems, with completely different minimax rates when compared to Huber’s setting. It is very unlikely that consistency towards $\theta^{*}$ is possible any longer in the realizable setting where ${\mathbb{P}_{X}}={P_{\theta^{*}}}$, and very likely that optimal estimators for Huber’s model behave extremely bad in this adversarial model. For instance, the average of extreme procedure becomes a terrible estimator under adversarial contamination of the mechanism: it is sufficient for the adversary to remove the $\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|-1$ smallest points and observe the largest one for a symmetric distribution to mislead the average of extremes estimator, no matter the value of the contamination ratio. At the opposite, the MMD estimator still behaves correctly in this setting, with an error which is (almost) linear in $\varepsilon$. The red curves in Figure 1 illustrate the behavior of the MLE, MMD and Average of Extremes Estimator: while the MMD is stable around $\varepsilon=0.1$, the average of extremes moves farther away from the true value of the mean as $n$ increases. The minimax optimality in the adversarial contamination setting is an open question.
@@
Figure 1: MLE, MMD (with Gaussian kernel) and Average of Extremes Estimator for a Gaussian mean estimation problem following Example 5.2 (Huber contamination) in light blue and Example 5.3 (Adversarial Contamination) in red. The shaded regions correspond to the empirical 50% quantile range obtained by rerunning the experiment 10001000 times for each $n$. In both settings, we simulate $X\sim N(0,1)$ without contamination. For the case of Huber contamination, the MCAR mechanism with $\alpha=0.5$ probability of missingness is contaminated by an MNAR mechanism for which $M=0$ iff $X>0$. For the adversarial mechanism, each $M_{i}$ is first generated following an MCAR mechanism with $\alpha=0.5$ probability of missingness, and in a second step, the $M_{i}$’s corresponding to the $2\varepsilon\cdot|\{i:M_{i}=0\}|-1$ smallest values are set to 1, while the $M_{i}$ corresponding to the largest value is set to 0. In both cases, $\varepsilon=0.1$ and $\epsilon=0$. Implementation was performed in R with the help of the R package regMMD (Alquier and Gerber, 2024a, ).
6 Simulation Study
In this section we empirically study the problem of Gaussian mean estimation under missing values and contamination. The base data distribution is $N(\mathbf{0},I_{d})$, the Gaussian distribution with identity covariance matrix. Following Chérief-Abdellatif and Alquier, (2022) we consider $n=500$, $d=10$ and contaminate the Gaussian distribution with a fraction $\epsilon=0.2$ of different distributions, in particular Gaussians with different means and delta measures, as shown in Table 1. As the baseline missing mechanism, we consider the following (blockwise) MCAR mechanism: With probability $\alpha_{m_{1}}$, $(X_{1},X_{2},X_{3})$ are missing, with probability $\alpha_{m_{2}}$, $(X_{3},X_{4},X_{5})$ are missing and with probability $\alpha_{m_{3}}$, $(X_{6},X_{7},X_{8})$ are missing. One could think of a measurement process, where the measurements itself are independent from each other, but if one fails, the others fail automatically. Finally with probability $\alpha_{m_{4}}$ all variables are observed. We choose $\alpha_{m_{j}}=0.25$ for all $j\in[4]$, so that around $1/4$ of draws are fully observed. We then contaminate this MCAR mechanism with a fraction $\varepsilon=0.2$ coming from a challenging self-censoring missingness mechanism. We adapt the self-censoring mechanism discussed in Miao et al., (2016): we define $\beta=d^{-1/2}\mathbf{1}=[1/\sqrt{d},\cdots,1/\sqrt{d}]^{\top}$ and
$\mathbb{Q}[M=m_{1}\mid X=x]$ $=\mathbb{Q}[M=m_{4}\mid X=x]=F(\beta^{\top}x)/2$
$\mathbb{Q}[M=m_{2}\mid X=x]$ $=\mathbb{Q}[M=m_{3}\mid X=x]=1/2-F(\beta^{\top}x)/2.$
Here $F$ is the cdf of $\beta^{\top}X$, where $X$ follows the contaminated distribution. For instance, if we contaminate with $N(\mu,I_{d})$, $F$ will be the cdf of a mixture of $\mathcal{N}(0,\beta^{\top}\beta)=\mathcal{N}(0,1)$ and $\mathcal{N}(\beta^{\top}\mu,1)$. This choice of parameters ensures that $\mathbb{Q}[M=m\mid X=x]$ is a valid distribution and that $\mathbb{Q}[M=m_{j}]=\alpha_{m_{j}}$ for $j\in\{1,\ldots,4\}$. As in Section 5.2, we then have ${\mathbb{P}[M=m|X=x]}=(1-\varepsilon)\alpha_{m}+\varepsilon{\mathbb{Q}[M=m|X=x]}$, with $\varepsilon=0.2$.
While mean estimation under contamination has attracted considerable research interest, less attention has been devoted to the setting with missing data. Notable recent exceptions include the approaches of Hu and Reingold, (2021); Ma et al., (2024), although these methods appear primarily designed for theoretical exploration rather than practical implementation. As a baseline, we thus include the coordinate-wise median as a simple but robust comparator. Under MCAR missingness, one can apply any robust method to the subset of fully observed data. We therefore incorporate the estimators “evfiltering” and “QUE” introduced in Diakonikolas et al., (2017, 2019) and Dong et al., (2019), respectively. Both methods rely on pruning observations using different criteria based on the assumed Gaussian distribution. We adopt the implementations from Anderson and Phillips, (2025), who adapted these algorithms for high-dimensional, small-sample settings. These methods have been shown to perform well under various contamination scenarios, particularly for $n=500$, $d=500$. Although we focus on $d=10$ here, both methods still perform strongly in the fully observed case. For our proposed MMD estimator, we use the Gaussian kernel with an implementation based on the regMMD R package (Alquier and Gerber, 2024a, ). The kernel bandwidth is selected via an adaptation of the median heuristic proposed by Gretton et al., 2012b .
Results for the root mean squared error (RMSE) are shown in Table 1. We note that in the first column, the Gaussian distribution is “contaminated” by the same distribution, corresponding to $\epsilon=0$. Thus, this setting corresponds to the reliasable contamination model in Ma et al., (2024), with contamination in the missing value mechanism only. Based on their results, we expect the sample mean to perform well in this setting. Indeed, the MLE performs extremely well here, though the MMD is not too far behind with the same performance as the median. However, as the degree of contamination is increased, the performance of the mean deteriorates rapidly, while the MMD estimate remains stable. The only exception to this is the case of $\delta_{\mathbf{1}}$, which, as discussed in Chérief-Abdellatif and Alquier, (2022), is a particularly bad contamination for MMD in this setting. Overall, the MMD is extremely competitive and in particular manages to beat the robust methods evfiltering and QUE. Although these methods are highly competitive, they only have access to around 1/4 of data points.
Compared to the similar simulation setting in Chérief-Abdellatif and Alquier, (2022) with complete data, we see that the RMSE is considerably higher throughout, indicating the increased difficulty of the problem. In particular, it appears that the effect of outliers on the mean worsens with missing data; while in Table 1 of Chérief-Abdellatif and Alquier, (2022) the RMSE of the mean is around 2 for the case of $N(\mathbf{10},I_{d})$ and $\delta_{\mathbf{10}}$, it is more than 3 times higher in Table 2. This is also the case without contamination of the MCAR mechanism, as shown in Appendix A.1. Remarkably, the RMSE values of the MMD in Table 1 (with MNAR contamination) and in Table 2 of Appendix A.1 (with $\varepsilon=0$) are nearly identical, suggesting that the additional MNAR contamination has little to no impact on the MMD estimator.
Table 1: RMSE under contamination of both the data distribution and missingness mechanism ($\varepsilon=\epsilon=0.2$). Parentheses indicate estimated standard deviations.
$N(\mathbf{0},I_{d})$ $N(\mathbf{1},I_{d})$ $N(\mathbf{10},I_{d})$ $\delta_{\mathbf{1}}$ $\delta_{\mathbf{10}}$
MMD 0.2 0.56 0.25 0.83 0.26
(0.06) (0.08) (0.06) (0.09) (0.05)
MLE 0.16 0.67 6.39 0.68 6.53
(0.03) (0.07) (0.65) (0.07) (0.64)
Median 0.2 0.64 1.06 1.08 1.11
(0.04) (0.08) (0.15) (0.13) (0.14)
evfiltering 0.3 0.86 0.36 0.89 0.37
(0.06) (0.15) (0.11) (0.13) (0.09)
QUE 0.3 0.86 3.92 0.89 5.28
(0.06) (0.15) (9.41) (0.13) (10.77)